OpenStudy (anonymous):

Which of the following is not an element of the following set: A: {positive integers divisible by 4}

OpenStudy (jim766):

cant see the following choices to answer it

OpenStudy (anonymous):

Thats all it says

OpenStudy (jim766):

if the choices have any negative numbers or zero, they wont be in it

OpenStudy (anonymous):

8 18 48 12

OpenStudy (jim766):

1,2,3,5,6,7,9,10,11,

OpenStudy (anonymous):

thats what i have to pick from

OpenStudy (jim766):

which one does not divide by 4 evenly

OpenStudy (jim766):

18/4 = ?

OpenStudy (jim766):

4.5

OpenStudy (anonymous):

thats not on there idk how to do this

OpenStudy (jim766):

in the question it tells you what is allowed to be in the set positive integers (numbers you count) and they also have to divide by 4 evenly all of the numbers they gave you are positive and divide evenly by 4 except 18 8/4 = 2 48/4 = 12 12/4 = 3

OpenStudy (jim766):

which is not an element of the set? 18

OpenStudy (anonymous):

thannksss